Object A of charge -2e-9 C is located on the x-axis at x = 0.4 m. Object B of charge 1e-9 C is located on the x-axis at x = 1 m. Point P of is located on the x-axis at x = 1.1 m. Determine the magnitude and direction of the net electric field point P by objects A and B.

861.265 N ⁄ C West
864.265 N ⁄ C East
863.265 N ⁄ C East
864.265 N ⁄ C West
863.265 N ⁄ C West
1
Expert's answer
2019-07-19T14:38:04-0400

The electric field at poit P

"E=E_1+E_2=k\frac{q_1}{AP^2}+k\frac{q_2}{BP^2}"

"=9\times 10^9\:\rm{N\cdot m^2/C^2}\left(\frac{-2\times 10^{-9}\:\rm{C}}{0.7^2\:\rm{m^2}}+\frac{1\times 10^{-9}\:\rm{C}}{0.1^2\:\rm{m^2}}\right)"

"=863.265\:\rm{N/C}\quad\rm{West}"